# Access Hoare Logic

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: XXXXdoi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: XXjournalnumber: Xarticle: 111publicationmonth: 8ccs: Theory of computation Hoare logicccs: Theory of computation Logic and verificationccs: Software and its engineering Software verificationccs: Security and privacy Logic and verificationccs: Security and privacy Software security engineering
@@ 1. Introduction
The correctness of programs has been a concern early on in the development of computers. The seminal contribution by Tony Hoare (Hoare, 1969) allows to reason about programs instruction by instruction using pre- and postconditions. It allows to ascertain that if the precondition is satisfied before execution, and program execution terminates, then the postcondition will be satisfied.
This form of ascertaining the correctness of programs has been and still is very successful. It is at the heart of several tools that are used in academia and industry to specify and verify software systems, such as SPARK (Adacore, 2025b) and Dafny (Dafny Development Team, 2025).
However, there are security properties that do not naturally fall within the framework of Hoare logic. Access control is the property of restricting access to a resource. It is a well-established area of computer security that includes identification, authorization, authentication, access approval, and auditing (Gollmann, 2011), where access approval is the process of deciding whether to grant or reject an access request. Access control has a broad range of applications, ranging from large distributed systems to local systems such as electronic keys. More recently, it gained importance in the context of blockchain and cryptocurrencies: Smart contracts are often written to control access to digital assets, such as cryptocurrencies, tokens, or other forms of digital assets managed by blockchain systems, thereby combining several properties of general access control. We will use the term access security to speak about such properties in general.
@@
Another form of reversing Hoare logic has been proposed in the literature (de Vries and Koutavas, 2011) and further developed to deal with incorrectness of programs (O’Hearn, 2019). That formalism expresses that for each final state $s^{\prime}$ satisfying the postcondition there is a state $s$ satisfying the precondition such that the program, when started in state $s$, can terminate in state $s^{\prime}$. This is a different perspective, which is incompatible with our approach, to the best of our knowledge.
@@
Alhabadi et al. (Alhabardi et al., 2022) have recently discovered that weakest preconditions for Hoare logic can be used to express access security for Bitcoin scripts. Our observation here expands on those insights and provides a reason why that is the case.
@@ 2.1. Access Security for Electronic Keys
Our first example is a digital key system in a hotel that uses cards to open doors, which was specified and verified by Jackson (Jackson, 2012) using Alloy, and by Nipkow (Nipkow, 2006) using Isabelle/HOL. In this system, the battery-powered door locks are not connected to a network. Instead, their state is controlled by cards that are inserted into them, or, in wireless systems, moved close to them. A lock holds an internal electronic key $\mathtt{dk}$ which controls its state. A card holds two electronic keys, $\mathtt{ck1}$ for the first key and $\mathtt{ck2}$ for the second. When a new guest receives a card (for example, from the hotel reception), it will contain the key corresponding to a previous guest as $\mathtt{ck1}$, and a new fresh key as $\mathtt{ck2}$. An access control program for the door lock suitable for use in such an electronic hotel key system should give access to the room in an access-secure way, by setting a variable $\mathtt{acc}$ to true: When the guest uses their card for the first time, access will be granted if the previous guest’s key is stored as the door key ($\mathtt{dk}{}=\mathtt{ck1}{}$), in which case the door key is updated to the new key $\mathtt{ck2}$, preventing the previous guest from gaining access to the room thereafter. The next time the new guest swipes their card, access will be granted because the second key on their card, $\mathtt{ck2}$, now matches the door key ($\mathtt{dk}{}=\mathtt{ck2}{}$).
@@ 2.2. Access Security for Bitcoin
In this context, the focus is on the access security of the locking script: In any successful run of the unlocking and then the locking script, the unlocking script is required to finish in a state which fulfills a suitable precondition, see Fig. 1. For a typical locking script like the standard locking script $\mathtt{scriptPubKey}_{\mathtt{pubKeyHash}}$ (Bitcoin Community, 2025), a typical precondition would express that an address and a signature are provided, where the address hashes to a value stored in the locking script, and the signature matches the address.
@@ 3. Access Hoare Triples
At this point we could introduce a calculus for access Hoare logic indirectly as negated Hoare logic in the sense of Theorem 3.1. However, it is of advantage to have a direct calculus for access Hoare logic. For example, a calculus for access Hoare logic would form the basis for systems that can verify access security in the sense of access Hoare logic using tools like theorem provers, similar to SPARK (Adacore, 2025b) and Dafny (Dafny Development Team, 2025) for verifying correctness based on standard Hoare logic. A reduction from access Hoare logic to standard Hoare logic, as outlined above, would introduce the contraposition of implications in verification proofs, i.e. $\neg B\rightarrow\neg A$ instead of $A\rightarrow B$. [^1] In more detail, a proof of an access Hoare triple $\langle P\rangle\,C\,\langle Q\rangle$ will usually involve subproofs of access Hoare triples of the form $\langle P^{\prime}\rangle\,C^{\prime}\,\langle Q^{\prime}\rangle$ and $\langle P^{\prime\prime}\rangle\,C^{\prime\prime}\,\langle Q^{\prime\prime}\rangle$, for certain subprograms $C^{\prime}$ and $C^{\prime\prime}$, where $C^{\prime}$ is immediately followed by $C^{\prime\prime}$. In such a situation, arguing access Hoare logic style from post- to preconditions requires proving the implication $P^{\prime\prime}\rightarrow Q^{\prime}$ from the precondition $P^{\prime\prime}$ of the second subproof to the postcondition $Q^{\prime}$ of the first. If we consider this situation by transformation to Hoare logic, we are faced with $\{\neg P\}\,C\,\{\neg Q\}$ and subproofs of $\{\neg P^{\prime}\}\,C^{\prime}\,\{\neg Q^{\prime}\}$ and $\{\neg P^{\prime\prime}\}\,C^{\prime\prime}\,\{\neg Q^{\prime\prime}\}$, now having to prove the implication $\neg Q^{\prime}\rightarrow\neg P^{\prime\prime}$ while arguing Hoare logic style from pre- to postconditions. For example, a straightforward implication such as $x=2\rightarrow x+2=4$ may translate into its contrapositive $\neg(x+2=4)\rightarrow\neg(x=2)$. While logically equivalent, [^2] In classical logic the latter is more difficult to understand and deal with in verification tasks. In general, this may even break verifiability. [^3] SPARK (Adacore, 2025b), which supports Hoare logic, tries to prove formulas generated by a verification process using an SMT solver. If that fails, SPARK offers to use the interactive theorem prover Coq/Rocq (Adacore, 2025a) instead. Another option when facing unprovability in SPARK, is to analyze formulas manually in Why3 (Why3, 2025; Adacore, 2025c), in order to find a reason for unprovability. In the last two cases, the syntactical form of formulas matters: For example, Coq/Rocq is based on intuitionistic logic, hence transformation of formulas need to respect intuitionistic logic.
@@ 4. The Calculus
We define the calculus for $\langle P\rangle\,C\,\langle Q\rangle$, which we denote access Hoare logic, following that for Hoare logic (Hoare, 1969), and show that it is sound with respect to the interpretation of access Hoare triples as given in (2) in the previous section. The development will focus on the $\mathtt{while}$ language similar to (Hoare, 1969), a standard imperative language. We will show completeness in Section 6.
@@ 6. Soundness and Completeness
Similar rules are known to be admissible for standard Hoare logic (Gordon, 2022). The soundness of both rules for access Hoare logic follows immediately. Their admissibility is also true: If the premises for a rule are provable in aHl, then they are true by soundness. But then the conclusion is also true by the soundness of the rule, which is therefore provable in aHl by completeness.
@@ 7. Conclusion
While access Hoare logic can be reduced to Hoare logic by introducing negation, we argued that this would have several negative consequences, including breaking verifiability using intuitionistic theorem provers — indeed, careful inspection shows that all proofs in this paper (except for Theorem 3.1, which needs to be replaced by Remark 3.2) only use intuitionistic logic, which we have confirmed using Agda (Agda Community, 2025) which is based on intuitionistic type theory. Details of that formalization will appear in a separate publication.
There are, of course, many areas that we did not touch upon. Basically, everything that has been investigated for Hoare logic should now be revisited in the context of access Hoare logic. Instead of aiming for a complete list, we mention two obvious areas. In our approach to soundness and completeness, we define the semantics of expressions and consequences in an informal way. There is a vast literature on ways to make this more precise, by considering formal semantics, their axiomatization and implications on Hoare logic (Apt, 1981). As a second area, the relationship to Dijkstra’s predicate transformer semantics and its variations need to be clarified (Dijkstra, 1975). Finally, detailed case studies are needed to demonstrate the usefulness of access Hoare logic for access security.
In connection to this paper, the authors have filed a patent entitled Verifying Access Security of a Computer Program (Beckmann and Setzer, 2025). This patent utilizes verification condition generation for access Hoare logic, which will be detailed in a separate publication.
@@ Acknowledgements.
We would like to thank Fahad F. Alhabardi. This article builds upon and expands the use of weakest preconditions for access security, which were explored during his PhD research (Alhabardi, 2024). We would also like to thank John V Tucker for valuable comments on an earlier draft of this paper.

### v2

@@ Abstract.
$c$s: Theory of computation Hoare logicccs: Theory of computation Logic and verificationccs: Software and its engineering Software verificationccs: Security and privacy Logic and verificationccs: Security and privacy Software security engineering
@@ 1. Introduction
The correctness of programs has been a concern early on in the development of computers. The seminal contribution by Tony Hoare (hoare:69) allows to reason about programs instruction by instruction using pre- and postconditions. It allows to ascertain that if the precondition is satisfied before execution, and program execution terminates, then the postcondition will be satisfied.
This form of ascertaining the correctness of programs has been and still is very successful. It is at the heart of several tools that are used in academia and industry to specify and verify software systems, such as SPARK (SPARKAda:Homepage) and Dafny (dafny:Homepage).
However, there are security properties that do not naturally fall within the framework of Hoare logic. Access control is the property of restricting access to a resource. It is a well-established area of computer security that includes identification, authorization, authentication, access approval, and auditing (gollmann:2011), where access approval is the process of deciding whether to grant or reject an access request. Access control has a broad range of applications, ranging from large distributed systems to local systems such as electronic keys. More recently, it gained importance in the context of blockchain and cryptocurrencies: Smart contracts are often written to control access to digital assets, such as cryptocurrencies, tokens, or other forms of digital assets managed by blockchain systems, thereby combining several properties of general access control. We will use the term access security to speak about such properties in general.
@@
Another form of reversing Hoare logic has been proposed in the literature (vriesKoutavas:ReverseHoareLogic) and further developed to deal with incorrectness of programs (OHearn:IncorrectnessLogic). That formalism expresses that for each final state $s^{\prime}$ satisfying the postcondition there is a state $s$ satisfying the precondition such that the program, when started in state $s$, can terminate in state $s^{\prime}$. This is a different perspective, which is incompatible with our approach, to the best of our knowledge.
@@
Alhabadi et al. (alhabadi:etal:2022) have recently discovered that weakest preconditions for Hoare logic can be used to express access security for Bitcoin scripts. Our observation here expands on those insights and provides a reason why that is the case.
@@ 2.1. Access Security for Electronic Keys
Our first example is a digital key system in a hotel that uses cards to open doors, which was specified and verified by Jackson (jackson2012:softwareAbstraction:2ndEd) using Alloy, and by Nipkow (Nipkow-VerficationHotelKeySystem) using Isabelle/HOL. In this system, the battery-powered door locks are not connected to a network. Instead, their state is controlled by cards that are inserted into them, or, in wireless systems, moved close to them. A lock holds an internal electronic key $\mathtt{dk}$ which controls its state. A card holds two electronic keys, $\mathtt{ck1}$ for the first key and $\mathtt{ck2}$ for the second. When a new guest receives a card (for example, from the hotel reception), it will contain the key corresponding to a previous guest as $\mathtt{ck1}$, and a new fresh key as $\mathtt{ck2}$. An access control program for the door lock suitable for use in such an electronic hotel key system should give access to the room in an access-secure way, by setting a variable $\mathtt{acc}$ to true: When the guest uses their card for the first time, access will be granted if the previous guest’s key is stored as the door key ($\mathtt{dk}{}=\mathtt{ck1}{}$), in which case the door key is updated to the new key $\mathtt{ck2}$, preventing the previous guest from gaining access to the room thereafter. The next time the new guest swipes their card, access will be granted because the second key on their card, $\mathtt{ck2}$, now matches the door key ($\mathtt{dk}{}=\mathtt{ck2}{}$).
@@ 2.2. Access Security for Bitcoin
In this context, the focus is on the access security of the locking script: In any successful run of the unlocking and then the locking script, the unlocking script is required to finish in a state which fulfills a suitable precondition, see Fig. 1. For a typical locking script like the standard locking script $\mathtt{scriptPubKey}_{\mathtt{pubKeyHash}}$ (bitcoin:wiki), a typical precondition would express that an address and a signature are provided, where the address hashes to a value stored in the locking script, and the signature matches the address.
@@ 3. Access Hoare Triples
At this point we could introduce a calculus for access Hoare logic indirectly as negated Hoare logic in the sense of Theorem 3.1. However, it is of advantage to have a direct calculus for access Hoare logic. For example, a calculus for access Hoare logic would form the basis for systems that can verify access security in the sense of access Hoare logic using tools like theorem provers, similar to SPARK (SPARKAda:Homepage) and Dafny (dafny:Homepage) for verifying correctness based on standard Hoare logic. A reduction from access Hoare logic to standard Hoare logic, as outlined above, would introduce the contraposition of implications in verification proofs, i.e. $\neg B\rightarrow\neg A$ instead of $A\rightarrow B$. [^1] In more detail, a proof of an access Hoare triple $\langle P\rangle\,C\,\langle Q\rangle$ will usually involve subproofs of access Hoare triples of the form $\langle P^{\prime}\rangle\,C^{\prime}\,\langle Q^{\prime}\rangle$ and $\langle P^{\prime\prime}\rangle\,C^{\prime\prime}\,\langle Q^{\prime\prime}\rangle$, for certain subprograms $C^{\prime}$ and $C^{\prime\prime}$, where $C^{\prime}$ is immediately followed by $C^{\prime\prime}$. In such a situation, arguing access Hoare logic style from post- to preconditions requires proving the implication $P^{\prime\prime}\rightarrow Q^{\prime}$ from the precondition $P^{\prime\prime}$ of the second subproof to the postcondition $Q^{\prime}$ of the first. If we consider this situation by transformation to Hoare logic, we are faced with $\{\neg P\}\,C\,\{\neg Q\}$ and subproofs of $\{\neg P^{\prime}\}\,C^{\prime}\,\{\neg Q^{\prime}\}$ and $\{\neg P^{\prime\prime}\}\,C^{\prime\prime}\,\{\neg Q^{\prime\prime}\}$, now having to prove the implication $\neg Q^{\prime}\rightarrow\neg P^{\prime\prime}$ while arguing Hoare logic style from pre- to postconditions. For example, a straightforward implication such as $x=2\rightarrow x+2=4$ may translate into its contrapositive $\neg(x+2=4)\rightarrow\neg(x=2)$. While logically equivalent, [^2] In classical logic the latter is more difficult to understand and deal with in verification tasks. In general, this may even break verifiability. [^3] SPARK (SPARKAda:Homepage), which supports Hoare logic, tries to prove formulas generated by a verification process using an SMT solver. If that fails, SPARK offers to use the interactive theorem prover Coq/Rocq (sparkAda:Doc:Coq) instead. Another option when facing unprovability in SPARK, is to analyze formulas manually in Why3 (why3:homepage; adacore:SPARKdoc:ArchitectureQualityAssurance:Why3), in order to find a reason for unprovability. In the last two cases, the syntactical form of formulas matters: For example, Coq/Rocq is based on intuitionistic logic, hence transformation of formulas need to respect intuitionistic logic.
@@ 4. The Calculus
We define the calculus for $\langle P\rangle\,C\,\langle Q\rangle$, which we denote access Hoare logic, following that for Hoare logic (hoare:69), and show that it is sound with respect to the interpretation of access Hoare triples as given in (2) in the previous section. The development will focus on the $\mathtt{while}$ language similar to (hoare:69), a standard imperative language. We will show completeness in Section 6.
@@ 6. Soundness and Completeness
Similar rules are known to be admissible for standard Hoare logic (gordon:LectureSlides:HoareLogic). The soundness of both rules for access Hoare logic follows immediately. Their admissibility is also true: If the premises for a rule are provable in aHl, then they are true by soundness. But then the conclusion is also true by the soundness of the rule, which is therefore provable in aHl by completeness.
@@ 7. Conclusion
While access Hoare logic can be reduced to Hoare logic by introducing negation, we argued that this would have several negative consequences, including breaking verifiability using intuitionistic theorem provers — indeed, careful inspection shows that all proofs in this paper (except for Theorem 3.1, which needs to be replaced by Remark 3.2) only use intuitionistic logic, which we have confirmed using Agda (agda:Documentation) which is based on intuitionistic type theory. Details of that formalization will appear in a separate publication.
There are, of course, many areas that we did not touch upon. Basically, everything that has been investigated for Hoare logic should now be revisited in the context of access Hoare logic. Instead of aiming for a complete list, we mention two obvious areas. In our approach to soundness and completeness, we define the semantics of expressions and consequences in an informal way. There is a vast literature on ways to make this more precise, by considering formal semantics, their axiomatization and implications on Hoare logic (apt:1981). As a second area, the relationship to Dijkstra’s predicate transformer semantics and its variations need to be clarified (dijkstra:1975). Finally, detailed case studies are needed to demonstrate the usefulness of access Hoare logic for access security.
In connection to this paper, the authors have filed a patent entitled Verifying Access Security of a Computer Program (beckmannSetzer:Patent:AccessHoareLogic:FilingUK). This patent utilizes verification condition generation for access Hoare logic, which will be detailed in a separate publication.
@@ Acknowledgements.
We would like to thank Fahad F. Alhabardi. This article builds upon and expands the use of weakest preconditions for access security, which were explored during his PhD research (alhabardi:PhD:SmartContractsAgda). We would also like to thank John V Tucker for valuable comments on an earlier draft of this paper.